# Finer: Investigating and Enhancing Fine-Grained Visual Concept Recognition in Large Vision Language Models

### v2

@@ 4.1 Probing for Concept-Related Parametric Knowledge
With the drastic performance degradation discussed in Section 3.2, we first need to verify whether concept-related knowledge already exists within the parametric knowledge space of the LVLMs to make sure that the models have already acquired concept-related knowledge necessary for zero-shot classification during the LLM pre-training stage. While we may assume that such knowledge has already been captured during pre-training of the LLM component on the Web-crawled concept documents, it is crucial to elucidate whether the knowledge exists within its parameters and whether the image representation is able to accurately capture such attributes of a concept that enables the models to infer the correct end concept. In this experiment, we measure the classification performance of the models with two different input types: (i) text-only input that consists of a concept’s external physical, descriptive attributes (details about the attribute extraction in §5), and (ii) image-only input that consists of the image of the concept; the generated outputs are measured in the same way as in Section 3.1. The text-only input, $\mathbf{X_{txt}}=[I;Attr;C]$, is composed of Instruction ($I$), web-extracted attributes ($Attr$), and coarse-grained labels ($C$); the output is the fine-grained concept label. Note, for fair comparison, we also include $C$ as input for image-only probing. In Figure 4, we show the results across three different datasets, iNaturalist, FGVC-Aircraft and Stanford Dogs. The results show that even with text-only input that contains the detailed physical attributes of a concept, LVLMs are capable of solving fine-grained visual classification, outperforming the image-only input, which is the model’s zero-shot performance on fine-grained classification. Additionally, we can also see that the larger the model size, the better the performance; this could be related to the amount of parametric knowledge that resides in a model’s LLM module.
@@ 4.2 Measuring the Modality Gap with Attribute Generation
Having observed the discrepancy between the visual and textual modalities, we now analyze whether LVLMs can observe and tell visually grounded physical attributes of an input image. We construct a set of Web-extracted concept attributes from Wikipedia documents (further elaborated in detail in Figure 5; §5) to be used as the reference texts. Then, we prompt the VLMs to generate a set of external, discriminative physical attributes of a concept when given either an image or text input (see the detailed prompts in Appendix 10); the text input refers to the concept label of the text along with a prompt that asks for the concept’s external physical descriptions. The textual similarities between the LVLM-generated attributes and the Web-extracted attributes are compared using five different scoring metrics that span both the token-level overlap and NLI model-based semantic similarity measure: ROUGE-1, 2, L (i.e., Longest-substring match), BertScore (Zhang et al., 2019), and AlignScore (Zha et al., 2023). Both the model-generated attributes and web-extracted attributes are linearized for textual similarity comparison; we detail the linearization process in Appendix A.2.
@@
As shown in Table 2, the text-only inputs show greater textual similarity to the Web-extracted reference attributes, indicating that while the attribute-specific knowledge acquired is being used by the textual modality, the visual modality does not leverage such knowledge to a degree that matches the textual modality. Aside from the modality discrepancy, a plausible interpretation is that these models fall short on accurately focusing on specific parts of the image, diluting away the fine-grained details such as stripes or patterns (Table 4) for a higher-level, semantic understanding of the image.
5 Finer: Fine-Grained Concept Categorization Benchmark
To facilitate the research for fine-grained image understanding in LVLMs, we propose a multiple granularity benchmark, Finer. Our benchmark intends to evaluate and improve the model capabilities in terms of improving the interplay between image understanding and textual latent knowledge of the concepts that appear within the images, based on the observation from previous sections (§3.2, 4)
@@ 5 Finer: Fine-Grained Concept Categorization Benchmark
Figure 5: Depiction of the Finer benchmark construction pipeline. Following the aggregation of the six benchmarks in the FGVC domain, concept attributes and concept images are retrieved from Wikipedia.
Dataset Name Total # of Instances (Test / Training) # of Superordinate Categories # of Coarse Categories # of Fine Categories # of Images Annotation Granularity Hierarchy Partial Attribute Annotation iNaturalist-2021 100K / 2.6M 11 1,103 10K 3,286,843 × CUB-200-2011 5,794 / 5,994 1 59 200 11,788 × FGVC-Aircraft 3,333 / 6,667 1 30 100 10,000 × Stanford Dogs 8,580 / 12,000 1 - 120 20,580 × × NABirds 24,633 / 2,929 1 146 404 48,562 Stanford Cars 8,144 / 8,041 1 7 196 16,185 × × FINER 372K/2.63M 16 1,416 11,171 3,393,958
@@ 5.1 Dataset Construction
We construct the benchmark based on six different FGVC datasets from the computer vision domain: iNaturalist-2021 (Van Horn et al., 2018), FGVC-Aircrafts (Maji et al., 2013), Stanford Dogs (Khosla et al., 2011), Stanford Cars (Krause et al., 2013), NABirds (Van Horn et al., 2015), and CUB-200-2011 (Wah et al., 2023). These datasets span a wide range of real-life objects such as airplanes, dogs, plants, birds, mammals and cars, challenging the models to cover a variety of fine-grained concepts.
@@
We first crawl all Wikipedia documents and their main images via a search API  [^2] https://github.com/goldsmith/Wikipedia in Python. The collected documents are then fed into GPT-4V (OpenAI, 2023) with an instruction to extract the document-grounded external, visible attributes of a concept. To briefly elaborate, we divide the extracted attributes into two different types: (i) required and (ii) likely attributes. The "required" attributes are the external, concept-indicative attributes that can be used for concept identification, e.g., blue-tailed hawks with black thorax with a broad apple green stripe, while the "likely" attributes are attributes that may co-occur with the concept, but may not be helpful in distinguishing the end concept, e.g., blue-tailed hawks inhabit trees; we provide the likely attributes as an additional meta-information for each concept. We also populate the dataset with superordinate and coarse-level concept labels. For example, as shown in Figure 5, we assign "Airplane" as the superordinate-level label and "Lockheed Martin" as the coarse-level label since FGVC-Aircrafts dataset provides a concept granularity hierarchy, e.g., Boeing $\rightarrow$ Boeing 707 $\rightarrow$ Boeing 707 MAX. However, datasets like Stanford Dogs do not provide neither the superordinate nor coarse-level labels for their fine-grained concepts. We therefore few-shot prompt GPT-4V to generate the coarse-grained labels and manually inspect their validity. We provide the dataset statistics in Table 3 and prompts for dataset generation in Appendix B.3, B.4.
@@
Image (Concept Name) Attributes (Text-only) Attributes (Image-only) Web-Extracted Attributes Dragonfly | Orthetrum Triangulare [Uncaptioned image] Blue tail ; Broad wings; White throat;Dark head; Banded tail Elongated body; Two pairs of wings; Six legs, Compound eyes , Segmented abdomen Dark face; Bluish eyes; Black thorax with a broad apple green stripe on both sides; Black segments 1-2 and 8-10 on the abdomen; Remaining segments of the abdomen pruinosed with azure blue Pinscher | Affenpinscher [Uncaptioned image] Black, gray, silver, or tan fur ; Rough, shaggy coat; Distinct "monkey-like" facial expression; Prominent chin and jaw; Small, round, dark eyes; Ears set high and usually cropped to a point Presence of fur; Four-legged stance; Distinct muzzle with a nose; Visible tail ; Ears that are either erect or floppy Harsh rough coat when not clipped; Shaggier coat over the head and shoulders forming a mane; Shorter coat over the back and hind quarters; Notable monkey-like expression; Coat is harsh and wiry in texture when properly maintained Embraer | Embraer ERJ 145 [Uncaptioned image] T-tail design ; Straight wing design with no winglets; Mounted engines on the rear fuselage; Small cockpit windows compared to the body size; Three sets of landing gear Fixed wings on either side of the fuselage ; Cockpit windows at the front of the fuselage; Jet engines, either under the wings or mounted on the rear of the fuselage; Landing gear with wheels for taking off T-tail configuration; Two rear-mounted Rolls-Royce AE 3007 series turbofan engines; Straight wing with no winglets; Narrow, tube-like fuselage; Short, nearly oval-shaped passenger windows.
@@
Table 4: Qualitative Analysis of the Text-only generated attributes and Image-only generated attributes from GPT-4V. The generated attributes based on the Text-only and Image-only dataset exhibits notable discrepancy. Juxtaposed with the Web-Extracted Attributes, Image-only attributes are less specific and explain more generic physical descriptions rather than those discriminative of the concept, while Text-only attributes capture finer-grained details of the concept. We provide the coarse-level and fine-level labels along with the images, and the bounding boxes are only drawn to match the highlighted text and are not part of the dataset.
@@ 5.2 Qualitative Analysis
Models CUB Stanford Dogs NABirds
LLaVA-1.5 2.744 30.990 2.078
LLaVA-1.5 (iNat) 2.709 33.473 1.562
LLaVA-1.5 (iNat-Attr) 5.177 32.354 5.728
Table 5: LLaVA-1.5 (7B) exhibits enhanced fine-grained performance after learning to generate textual attributes based on image input. (iNat) is fine-tuned directly on the concept label and (iNat-Attr) is fine-tuned to generate the attributes first and then predict the concept label.
@@
In Table 4, we provide the Web-extracted attributes from our dataset, along with model-generated attributes as a case study on the modality gap phenomenon from Section 4. The attributes are from GPT-4V, but note that the generated attributes from other models such as LLaVA-1.5, all exhibit a similar trend. When provided with image-only input, the model generates a set of attributes that pertain to the image, e.g., for dragonfly it describes its elongated body and two pairs of wings. However, the attributes from image-only input are more generic compared to those from text-only input. In other words, these attributes are not discriminative enough to identify the input image as a specific concept category. This again suggests that these models not only fail to properly observe the fine-grained details of a concept, but fail to leverage the knowledge contained within its own parameters as can be seen from the outputs of text-only inputs.
@@
5.3 Zero-Shot Transferability of Learning to Generate Attributes
To substantiate the effectiveness of the Web-extracted attributes in Finer, we construct an instruction-tuning mixture with the AttrSeek pipeline using the iNaturalist subset on top of LLaVA-1.5’s instruction-tuning dataset (see Appendix A.4). We fine-tune LLaVA-1.5 (7B) and evaluate the fine-tuned model on three different held-out datasets. To briefly elaborate, each instance in our mixture follows three steps: (i) Ask the model to identify the course-grained concept category, (ii) generate the visual attributes of the concept, (iii) predict the fine-grained concept label. In Table 5, we see that the model outperforms simple fine-tuning and zero-shot classification performance on the three held-out datasets. This result suggests that in FGVC, instruction-tuning LVLMs by first forcing the model to look at the apparent external, visible attributes in the image by explicitly generating the textual attributes, and then subsequently performing classification improves the model’s performance. For additional details on the instruction-tuning on our mixture, refer to Appendix A.
@@ 6 Discussion and Conclusion
In this paper, we investigate the substantial lack of fine-grained visual comprehension (FGVC) ability of instruction-tuned LVLMs. Our empirical results reveal that significant modality gap exists between the visual and textual modalities, preventing the models from effectively leveraging textual knowledge for zero-shot fine-grained visual classification. Furthermore, our results also demonstrate that while LVLMs are able to describe generic attributes of a concept, they exhibit hallucination and are unable to ground their explanation on the concept present within an input image. To establish zero-shot FGVC research among LVLMs and facilitate their in-depth understanding of apparent attributes, we proposed a novel multiple granularity, attribute-centric benchmark, Finer.
@@ A.1 Hyperparameter Settings of Vision-Language Models
Table 6: Hyperparameters of the Vision-Language Models (VLMs) studied in this work. LLaVA refers to LLaVA-1.5 and InstBLIP refers to the InstructBLIP model. All the VLMs use the same hyperparameter settings for both the 7B and 13B variants; the same applies to Flan-T5-XL and Flan-T5-XXL models.

### v4

@@ 4.1 Probing for Concept-Related Parametric Knowledge
With the drastic performance drop in Section 3.2, we first need to verify whether concept-attribute knowledge already exists within the LVLM parameters to make sure that the models have already acquired the knowledge necessary for zero-shot classification. The concept-attribute knowledge refers to the textual parametric knowledge related to specific concepts, e.g., a concept Bengal Tiger has visual attributes dark brown or black stripes.
In this experiment, we measure the classification performance of the models with two different input types: (i) Text-only input that consists of a concept’s external visual attributes (details about the attribute extraction in §5), and (ii) Image-only input that consists of the image of the concept. The text-only input, $\mathbf{X_{txt}}=[I;Attr;C]$, is composed of Instruction ($I$), visual attributes ($Attr$), and coarse-grained labels ($C$); the image-only input is $\mathbf{X_{v}}=[I;X_{img};C]$. The final output is the concept name. Note, for fair comparison, we also include $C$ as input for image-only probing. Refer to Appendix B.2 for prompts. In Figure 4, the results show that even with text-only input that contains the detailed physical attributes of a concept, LVLMs are capable of solving fine-grained visual classification, outperforming the image-only input. The results imply two things: (i) concept-attribute knowledge exists in model parameters, and (ii) while the visual attributes are strongly correlated with the concept, the image modalities are incapable of leveraging the concept-attribute knowledge. We also see that the larger the model size, the better the performance, relating to the amount of parametric knowledge that resides in the LLMs.
@@ 4.2 Measuring the Modality Gap with Attribute Generation
Having observed the discrepancy between the visual and textual modalities, we now analyze whether LVLMs can observe and tell visually grounded physical attributes of an input image. We construct a set of Web-extracted concept attributes from Wikipedia documents (further elaborated in detail in Figure 6; §5) to be used as the reference texts. Then, we prompt the models to generate a set of external, discriminative physical attributes of a concept when given either an image or text input (see the detailed prompts in Appendix 11); the text input refers to the concept label along with a prompt that asks for the concept’s visual attributes. The textual similarities between the LVLM-generated attributes and the web-extracted attributes are compared using five different scoring metrics that span both the token-level overlap and NLI model-based semantic similarity measure: ROUGE-1, 2, L, BertScore (Zhang et al., 2019), and AlignScore (Zha et al., 2023). Both the model-generated attributes and web-extracted attributes are linearized for textual similarity comparison (Appendix A.2).
@@
As shown in Table 2, text-only inputs show greater textual similarity to reference attributes, indicating that while the concept-attribute knowledge is being used by the textual modality, the visual modality does not leverage such knowledge to a degree that matches the textual modality. Aside from the modality discrepancy, these models potentially fall short on accurately focusing on specific parts of the image, diluting away the fine-grained details such as stripes or patterns (Table 3) for a coarse-grained understanding of the image.
4.3 Visual Information Loss After Projection
Figure 5: Linear Probing on Projected Image Embeddings. Classification accuracy (%) for before and after image embedding projection to textual space.
We also evaluate the impact of projection from the visual embedding space to the textual space through linear probing. We use CLIP-ViT-L/14 as the image encoder and use LLaVA-1.5’s projector. We freeze both the image encoder and the projector and finetune a multi-layer perceptron (MLP) layer on top for classification for 10 epochs (for experiment details refer to Appendix B.5). As shown in Figure 5, the loss of visual information encoded by the vision encoder leads to substantial drop in classification performance across the six FGVC tasks. The results strongly suggest that such loss of visual information further contributes to the modality gap between the two modalities, especially when performing tasks such as FGVC that rely heavily on fine-grained visual attributes.
@@ 4.3 Visual Information Loss After Projection
Figure 6: Depiction of the Finer benchmark construction pipeline. Following the aggregation of the six benchmarks in the FGVC domain, concept attributes and concept images are retrieved and extracted from Wikipedia documents.
5 Finer: Fine-Grained Concept Recognition Benchmark
To facilitate research for fine-grained image understanding in LVLMs, we propose a new benchmark and training mixture, Finer. Finer intends to evaluate the interplay between the concept image understanding and attribute knowledge, in addition to the training mixture that mitigates the modality gap and enhances fine-grained concept recognition.
@@ 5.1 Dataset Construction
We construct Finer based on six different FGVC datasets: iNaturalist-2021 (Van Horn et al., 2018), FGVC-Aircrafts (Maji et al., 2013), Stanford Dogs (Khosla et al., 2011), Stanford Cars (Krause et al., 2013), NABirds (Van Horn et al., 2015), and CUB-200-2011 (Wah et al., 2023). These datasets span a wide range of objects such as airplanes, insects, plants, birds, mammals and cars, challenging the models to cover a variety of fine-grained concepts.
@@
We first crawl all Wikipedia documents and their main images via a search API [^1] https://github.com/goldsmith/Wikipedia. We then extract external, visual attributes of a concept (i.e., concept-indicative attributes) with GPT-4V as our attribute extractor (OpenAI, 2023) for its strong zero-shot text span extraction ability (Huang et al., 2024). To briefly elaborate, we divide the extracted attributes into two different types: (i) required and (ii) likely attributes. The "required" attributes are the external, concept-indicative attributes that can be used for concept identification, e.g., blue-tailed hawks with black thorax with a broad apple green stripe, while the "likely" attributes are attributes that may co-occur with the concept but is not directly correlated with the concept, e.g., blue-tailed hawks inhabit trees; we provide the likely attributes as meta-information since existing models such as MetaFormer (Diao et al., 2022) has proven that meta-information associated with these fine-grained concepts are beneficial for more accurate FGVC performance; however, since the use of meta-information for better FGVC is not the main focus of this work, we leave it to future works to leverage this information. We also populate the dataset with superordinate and coarse-level concept labels for multi-granular concept recognition performance evaluation. For example, as shown in Figure 6, we assign Airplane as the superordinate-level label and Lockheed Martin as the coarse-level label since FGVC-Aircrafts dataset provides a concept granularity hierarchy, e.g., Boeing $\rightarrow$ Boeing 707 $\rightarrow$ Boeing 707 MAX. However, datasets like Stanford Dogs do not provide such granular hierarchy for their fine-grained concepts. We therefore few-shot prompt GPT-4V to generate the coarse-level labels and manually inspect their validity. We provide the dataset statistics in Table 9 and extraction prompts in Appendix B.3 and B.4.
@@
Image (Concept Names) Attributes (Text-only) Attributes (Image-only) Finer Attributes Dragonfly | Orthetrum Triangulare [Uncaptioned image] Blue tail ; Broad wings; White throat;Dark head; Banded tail Elongated body; Two pairs of wings; Six legs, Compound eyes , Segmented abdomen Dark face; Bluish eyes; Black thorax with a broad apple green stripe on both sides; Black segments 1-2 and 8-10 on the abdomen; Remaining segments of the abdomen pruinosed with azure blue Pinscher | Affenpinscher [Uncaptioned image] Black , gray, silver, or tan fur; Rough, shaggy coat; Distinct "monkey-like" facial expression; Prominent chin and jaw; Small, round, dark eyes; Ears set high and usually cropped to a point Presence of fur; Four-legged stance; Distinct muzzle with a nose; Visible tail ; Ears that are either erect or floppy Harsh rough coat when not clipped; Shaggier coat over the head and shoulders forming a mane; Shorter coat over the back and hind quarters; Notable monkey-like expression; Coat is harsh and wiry in texture when properly maintained Embraer | Embraer ERJ 145 [Uncaptioned image] T-tail design ; Straight wing design with no winglets; Mounted engines on the rear fuselage; Small cockpit windows compared to the body size; Three sets of landing gear Fixed wings on either side of the fuselage ; Cockpit windows at the front of the fuselage; Jet engines, either under the wings or mounted on the rear of the fuselage; Landing gear with wheels for taking off T-tail configuration; Two rear-mounted Rolls-Royce AE 3007 series turbofan engines; Straight wing with no winglets; Narrow, tube-like fuselage; Short, nearly oval-shaped passenger windows.
@@
Table 3: Qualitative Analysis of the Text-only and Image-only generated attributes from GPT-4V against Finer. The generated attributes based on the Text-only and Image-only dataset exhibits notable discrepancy. Juxtaposed with our Finer Attributes, Image-only attributes are not concept-indicative and generic compared to Text-only which are discriminative of the concept. We provide Finer attributes as a reference for comparison. We provide the coarse-level and fine-level labels along with the images, and the bounding boxes are only drawn to match the highlighted text and are not part of the dataset.
@@ 5.2 Qualitative Analysis
In Table 3, we provide model-generated attributes as a case study on lack of visually-grounded generation. The attributes are from GPT-4V, but note that the generated attributes from other models, such as LLaVA-1.5, all exhibit a similar trend. When provided with image-only input, the model generates a set of attributes that pertain to the image, e.g., elongated body and two pairs of wings for dragonfly. However, the attributes from image-only input are non-discriminative compared to those from text-only input; furthermore, changing the prompting technique to elicit more fine-grained, detailed physical attributes from the LLMs either lead to hallucination or needlessly verbose outputs that describe non-concept related aspects of the input image. In other words, these attributes do not serve as useful knowledge to identify the input image as a specific concept. This again suggests that these models not only fail to properly observe the fine-grained details of a concept, but fail to leverage the knowledge contained within its own parameters as can be seen from the outputs of text-only inputs.
@@
5.3 Enhanced Zero-Shot Transferability from Learning to Generate Attributes
Models CUB-200-2011 Stanford Dogs NABirds iNaturalist Stanford Cars FGVC-Aircrafts
LLaVA-1.5 (7B) 5.262 16.718 3.145 6.788 25.317 29.314
Direct Prediction 21.071 22.942 12.610 7.109 24.624 28.622
Finer 20.673 36.297 13.692 7.530 29.974 32.293
Table 4: Zero-shot Performance on FGVC. Finetuning on the Finer mixture significantly enhances the zero-shot performance on all six FGVC tasks. We choose LLaVA-1.5 (7B) for this experiment with 1 dataset held-out and 5 other datasets held-in for finetuning. Direct Prediction refers to the setting without the AttrSeek pipeline and the model directly predicting the final concept label without the intermediate attribute generation
To substantiate the effectiveness of our visual attributes in Finer, we construct an instruction-tuning mixture based on the AttrSeek prompting pipeline to improve the zero-shot attribute generation and FGVC performance (see Appendix A.4). The Finer mixture consists of six subsets, where each split is a 5 held-in and 1 held-out FGVC datasets for training and evaluation, respectively. For instance, for the iNaturalist subset, the iNaturalist set is not included in training for zero-shot evaluation. Each instance of the training mixture follows the AttrSeek pipeline (§3.2). In Table 4, we finetune LLaVA-1.5 (7B) on the training mixture and see that the Finer-tuned model outperforms the direct finetuned counterpart that was simply trained to directly predict the concept label. This implies that in FGVC, instruction-tuning LVLMs to attend to visible attributes in images by explicitly generating the attributes and then subsequently performing classification improves the model’s performance. Our interpretation is that the generation of the attributes of concepts in the images allow the model to leverage its concept-attribute parametric knowledge (identified in Figure 4), and perform better zero-shot FGVC classification. It also underscores the effectiveness of the AttrSeek pipeline in model training to improve the inherent, zero-shot capability of LVLMs for fine-grained concept recognition. We demonstrate case studies in Table 5, where we present the zero-shot generated outputs of the FINER mixture-trained LLaVA-1.5 (7B) and the one that was only finetuned to directly predict the final concept label. Refer to Appendix A for additional details on training.
Image (Concept Names) Attributes from Finer Attributes from Direct Pred. Warbler | Yellow-headed Warbler [Uncaptioned image] It has yellow round head, white body with a hint of gray, black eyes and black-colored tail. In the image, the bird is perched on a wire, and it has a yellow beak. Its physical characteristics include a white body, a black tail, and a yellow beak. Albatross | Black-footed Albatross [Uncaptioned image] Almost all black plumage, White markings around the base of the beak, White markings below the eye, Dark beak and feet , White undertail coverts in some adults. The bird has Predominantly dark color with subtle variations between brown and gray. The face appears to have lighter shades of brown or gray.
Table 5: Qualitative analysis of the zero-shot outputs generated by LLaVA-1.5 (7B) instruction-tuned on Finer and Direct Prediction dataset. The generated attributes from the Finer trained LLaVA-1.5 (7B) generates accurate, image-grounded outputs when compared to the Direct Prediction counterpart. The attributes are indicated by their discriminative characteristics in contrast to more generic/hallucinated ones. We provide the coarse-level and fine-level labels along with the images, and the bounding boxes are only drawn to match the highlighted text and are not part of the dataset.
@@ 6 Discussion and Conclusion
In this paper, we provide an in-depth analysis on the substantial lack of fine-grained visual comprehension (FGVC) ability among instruction-tuned LVLMs. Our work discovers the presence of modality gap in LVLMs and propose a prompting scheme, AttrSeek, and training mixture, Finer to mitigate the gap and improve the zero-shot FGVC ability of LVLMs. We also reveal that the loss of visual information after projection hinder the effective cross-modal interplay from manifesting. Such discrepancy leads to LVLMs being unable to exploit the rich parametric knowledge and deteriorates performance in visual concept recognition. While this is the first study of FGVC among instruction-tuned LVLMs, we hope our work would further the research endeavors in this direction.
@@ A.1 Hyperparameter Settings of Vision-Language Models
Table 6: Hyperparameters of the Vision-Language Models (VLMs) studied in this work. LLaVA refers to LLaVA-1.5 and InstBLIP refers to the InstructBLIP model. All the VLMs use the same hyperparameter settings for both the 7B and 13B variants; the same applies to Flan-T5-XL and Flan-T5-XXL models, which consist of 3B and 11B parameter, respectively.